OpenStudy (anonymous):

The quotient of the exact division between two numbers a and b is equal to the gcf(a,b). If the sum of the gcf(a,b) and the lcm(a,b) is equal to 56, find ab.

OpenStudy (ddcamp):

413: This was admittedly sort of guess and check, but I figured if a/b is the gcf(a,b) a is probably the square root of b. The lcm would then be b. Plugging in for a = 7, you get the GCF to be 7 and the lcm to be 49. 49+7 = 56. A = 49, B = 7 A*B = 413

OpenStudy (anonymous):

Yes!49 and 7 are correct!, By the way 49*7=343. Thanks!.

OpenStudy (ddcamp):

I have no idea how I got 413. Silly calculators.